ADNAN MUHAMMED

Grade 12,

A 20N metal block is suspended by a spring balance. A beaker containing some water is placed on the weighing machine reads 36N. The spring balance is now lowered so that the block gets immersed in water. The spring balance now reads 16N. What will be the reading of the weighing machine ?


 A 20N metal block is suspended by a spring balance. A beaker containing some water is placed on the weighing machine reads 36N. The spring balance is now lowered so that the block gets immersed in water. The spring balance now reads 16N. What will be the reading of the weighing machine ?

Grade:

2 Answers

Aman Bansal
13 years ago

dear kumar gaurav,

The spring balance reading decreases by (20-16) 4N when the block is immersed in water,

therefore the upthrust on the block is 4N.

according to the third law of motion it will exert an equal and opposite force

therefore the reading of the weighing machine will be

36+4 = 40 N

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL

Agniva Bagchi
6 years ago
initial reading =20N 
Final reading =16N
decreased weight=20-16=4N(upward direction)
now weight recorded by weighing machine=36N
therefore the reading of the weighing machine will be 36+4=40N(downward direction)(acc.to newton’s third law)
 
if you like the give a thums up!!!